# Saliency-Guided Training for Fingerprint Presentation Attack Detection

### v1

@@ 1 Introduction
Alongside contributions at the intersection of saliency-guided training and fingerprint PAD, this paper releases111Instructions on how to request a copy of the dataset will be released upon publication of the peer-reviewed version. all explored saliency data, including the 800 doubly human-annotated saliency maps, the various sets of algorithmically-sourced pseudosaliency and the trained fingerprints PAD models.
@@ 3.1 Justification of Data Selection
We use the LivDet-Fingerprint datasets from 2015, 2017, 2019, and 2021 for both saliency collection and saliency-guided training [20, 21, 25, 5]. (The LivDet-Fingerprint 2023 dataset was not available to us at the time of writing this paper.) These datasets were chosen for three key reasons. First, they are the most recent and well-established benchmarks for evaluation of fingerprint PAD available to these authors. Second, their combined use provides a broader range of presentation attack instruments and capture devices, providing greater diversity than any single LivDet-Fingerprint competition alone. Finally, the LivDet-Fingerprint 2021 test set is the most challenging among our available datasets, as evidenced by its lower reported competitor scores compared to previous available LivDet-Fingerprint test sets. This increased difficulty, along with its inclusion of attack types not present in its training set or any prior LivDet dataset, makes it a strong choice for evaluating model performance.
@@ 3.2 Construction of Limited Data Context
The training of biometric PAD models, including fingerprint PAD, is often limited by the availability of high-quality spoof samples [17]. To simulate training conditions under a limited data availability context, we construct an 800-sample training dataset. The limited data context is balanced by class, with 400 bonafide and 400 spoof samples. The spoof samples are further balanced over eight attack types at 50 samples each, sourced from the LivDet-Fingerprint 2015-2019 train and test datasets and the LivDet-Fingerprint 2021 train datasets: Ecoflex [20, 21, 25], gelatine [20, 21, 25], latex [20, 21, 25, 5], liquid Ecoflex [20, 21, 25], woodglue [20, 21, 25], body double [21], mix [25], and RPRO Fast [5]. Bonafide and spoof samples are further balanced by the capture sensors represented by each subclass (note that not all attack types are represented by all scanners).
@@ 3.3 Human-Annotative Saliency Collection
We recruited 50 participants from the University of Notre Dame using a convenience sampling technique due to accessibility, scheduling, and other practical constraints. We collected annotations over the limited 800-sample dataset described in Sec. 3.2. Each participant annotated 32 fingerprint scans, 16 bonafide and 16 spoof samples, with spoof samples spanning eight attack types (two images per type). Participants used an image annotation software specially designed for this study to (a) predict each sample’s class (genuine or fake), (b) annotate salient regions supporting their decision, and (c) optionally provide textual description to further support their classification decisions (e.g., in cases when it would be difficult to support subject’s decision by annotating the image, such as perfect symmetry of the fingerprint image, or unnatural contrast, or too dark image). The created image annotation software also recorded start and end time stamps of each stroke. This allows for fine-grained analysis of annotation time and provides the order of features annotated by each subject on each sample. Confounding factors in this study include the level of biometric expertise or fingerprint familiarity among annotators. However, we consider this factor to be of limited concern, as prior saliency collection studies have demonstrated success with non-expert annotators [2, 3].
@@ 5 Results
In limited-data loss-based guidance (scenario S3), the human-guided configuration achieving the highest accuracy is configured with the Area of Interest saliency granularity, the Inception architecture, and a CYBORG weight of 0.9. Its achieved accuracy of 0.879±0.006 and AUC of 0.959±0.002 outperform the best baseline configuration (scenario S2) by 0.017 and 0.013, respectively.
@@
The generative nature of pseudosaliency enables saliency-guided training to scale effectively in larger data availability contexts, as explored in scenario S5. The best configuration in this scenario employs minutiae-based Area of Interest pseudosaliency, achieving an accuracy of 0.938±0.004. This result outperforms the large-data baseline (scenario S1) by 0.027. When compared to the reported LivDet-Fingerprint 2021 rankings [5], this performance would have earned or tied for first place, accounting for our reported standard deviation (which was not included in the LivDet-Fingerprint 2021 results). This configuration raises spoof detection accuracy by 8.9%, demonstrating the generalization strength of saliency-guided training. It is particularly notable given that all spoof samples in the LivDet-Fingerprint 2021 test set represent unseen attack types.
@@ 6 Discussion
A closer look at model performance metrics in Tab. 1 reveals that saliency-guided training primarily raises spoof detection accuracy, with bonafide accuracy usually affected to a lesser degree. In the limited-data context, the best configuration (Inception, CYBORG, =0.7, Minutiae FOI) improves spoof accuracy by 4.3%, while bonafide accuracy increases by only 0.4%. Similarly, in the large-data context, the best configuration (DenseNet, CYBORG, =0.9, Minutiae AOI) gains 8.9% in spoof accuracy but loses 3.1% in bonafide accuracy. As mentioned in Sec. 5, this behavior is particularly noteworthy considering that spoof samples in the LivDet-Fingerprint 2021 test set are entirely novel. This improvement in spoof performance highlights the potential of saliency-guided training to improve model generalization.
@@ 6.3 Blurring and BOI Saliency Granularity
In blur-based saliency guidance, the Boundary of Interest (BOI) saliency granularity consistently performs well. It is the best granularity for human saliency and second best among explored pseudosaliency types. BOI saliency identifies a minimal enclosing rectangle around all regions of interest, which typically encompasses most of the fingerprint area. Consequently, when non-salient regions are blurred, the fingerprint pattern remains largely unaffected. Combined with eightfold data augmentation through multiple blur radii, the model encounters the nearly unblurred pattern repeatedly. This inadvertently increases exposure to samples during training, similar to training with standard cross entropy for many additional epochs. While BOI saliency achieves high scores on key metrics, this effectiveness is likely due to unintended augmentation behavior rather than true saliency-based guidance. Its success raises questions about whether performance gains come from saliency-guided learning or from increased data redundancy, which impacts how blur-based saliency guidance results are interpreted.
@@ 7 Conclusion
Accuracy: Classification accuracy on the LivDet-Fingerprint 2021 test set.
Live Acc.: Classification accuracy on live (bonafide, real) fingerprint samples from the LivDet-Fingerprint 2021 test set.
Spoof Acc.: Classification accuracy on spoof (fake) fingerprint samples from the LivDet-Fingerprint 2021 test set.
@@
Appendix A Scenario S1 Results
| Network | AUC | Accuracy | Live Acc. | Spoof Acc. | Placement | d’ | FNR@FPR=1% |
| ResNet DenseNet Inception | 0.989±0.000 0.990±0.001 0.990±0.010 | 0.911±0.001 0.905±0.003 0.911±0.008 | 0.995±0.000 0.995±0.001 0.996±0.001 | 0.821±0.003 0.808±0.006 0.819±0.018 | 4 4 3-4 | 3.261±0.020 3.317±0.045 3.273±0.072 | 0.133±0.004 0.146±0.010 0.137±0.006 |
@@ Appendix B Scenario S2, Full Results
| Architecture | AUC | Accuracy | Live Acc. | Spoof Acc. | Placement | d’ | FNR@FPR=1% |
| ResNet DenseNet Inception | 0.930±0.003 0.918±0.016 0.946±0.007 | 0.847±0.003 0.839±0.012 0.862±0.010 | 0.928±0.007 0.944±0.004 0.948±0.009 | 0.760±0.014 0.725±0.025 0.769±0.014 | 5 5 5 | 2.093±0.027 1.978±0.161 2.283±0.094 | 0.525±0.042 0.518±0.063 0.474±0.027 |
@@ Appendix C Scenario S3, Full Results
| Network | Alpha () | AUC | Accuracy | Live Acc. | Spoof Acc. | Placement | d’ | FNR@FPR=1% |
| ResNet DenseNet Inception | 0.9 | 0.936±0.006 0.944±0.005 0.953±0.003 | 0.860±0.012 0.861±0.009 0.872±0.003 | 0.934±0.011 0.952±0.008 0.957±0.003 | 0.779±0.035 0.763±0.012 0.780±0.004 | 4-5 5 4-5 | 2.153±0.068 2.254±0.057 2.365±0.037 | 0.510±0.043 0.448±0.039 0.446±0.008 |
| ResNet DenseNet Inception | 0.7 | 0.932±0.006 0.936±0.004 0.952±0.003 | 0.850±0.010 0.847±0.002 0.872±0.003 | 0.942±0.010 0.943±0.007 0.952±0.004 | 0.751±0.022 0.743±0.008 0.785±0.010 | 5 5 4-5 | 2.116±0.063 2.149±0.050 2.360±0.041 | 0.526±0.034 0.512±0.027 0.482±0.061 |
| ResNet DenseNet Inception | 0.5 | 0.920±0.005 0.933±0.006 0.954±0.005 | 0.835±0.005 0.852±0.010 0.864±0.014 | 0.928±0.007 0.938±0.013 0.954±0.007 | 0.735±0.017 0.760±0.028 0.767±0.037 | 5 5 4-5 | 1.992±0.048 2.116±0.064 2.387±0.070 | 0.585±0.027 0.566±0.045 0.480±0.019 |
| ResNet DenseNet Inception | 0.3 | 0.901±0.006 0.915±0.010 0.945±0.004 | 0.820±0.004 0.830±0.014 0.861±0.006 | 0.902±0.013 0.927±0.001 0.938±0.007 | 0.731±0.009 0.724±0.029 0.778±0.020 | 5-7 5-7 5 | 1.823±0.051 1.945±0.093 2.267±0.050 | 0.689±0.045 0.596±0.011 0.555±0.056 |
| ResNet DenseNet Inception | 0.1 | 0.862±0.012 0.872±0.014 0.926±0.006 | 0.776±0.013 0.787±0.014 0.844±0.007 | 0.878±0.013 0.889±0.010 0.899±0.008 | 0.665±0.016 0.677±0.028 0.784±0.011 | 7-10 7-9 5 | 1.540±0.079 1.607±0.093 2.044±0.056 | 0.839±0.057 0.762±0.029 0.671±0.030 |
| Network | Alpha () | AUC | Accuracy | Live Acc. | Spoof Acc. | Placement | d’ | FNR@FPR=1% |
| ResNet DenseNet Inception | 0.9 | 0.908±0.009 0.931±0.013 0.959±0.002 | 0.819±0.002 0.838±0.022 0.879±0.006 | 0.927±0.009 0.949±0.013 0.966±0.002 | 0.703±0.006 0.717±0.050 0.786±0.012 | 6 5-7 4-5 | 1.881±0.074 2.111±0.145 2.462±0.036 | 0.595±0.024 0.511±0.046 0.379±0.039 |
| ResNet DenseNet Inception | 0.7 | 0.921±0.017 0.925±0.006 0.953±0.008 | 0.838±0.022 0.833±0.003 0.865±0.013 | 0.926±0.016 0.946±0.007 0.957±0.007 | 0.743±0.037 0.712±0.014 0.766±0.022 | 5-7 5 4-5 | 2.005±0.159 2.038±0.060 2.377±0.116 | 0.556±0.095 0.553±0.013 0.424±0.029 |
| ResNet DenseNet Inception | 0.5 | 0.927±0.006 0.932±0.009 0.941±0.002 | 0.840±0.007 0.840±0.014 0.857±0.002 | 0.940±0.008 0.944±0.005 0.942±0.005 | 0.731±0.014 0.727±0.026 0.767±0.003 | 5 5 5 | 2.056±0.059 2.114±0.101 2.207±0.024 | 0.581±0.051 0.562±0.034 0.552±0.021 |
| ResNet DenseNet Inception | 0.3 | 0.892±0.011 0.911±0.014 0.943±0.009 | 0.811±0.011 0.823±0.010 0.859±0.013 | 0.911±0.011 0.939±0.014 0.937±0.004 | 0.704±0.015 0.699±0.006 0.776±0.024 | 5-7 5-7 4-5 | 1.754±0.085 1.912±0.121 2.246±0.108 | 0.705±0.052 0.622±0.035 0.549±0.047 |
| ResNet DenseNet Inception | 0.1 | 0.854±0.004 0.862±0.019 0.915±0.010 | 0.780±0.003 0.782±0.020 0.836±0.012 | 0.848±0.008 0.880±0.007 0.891±0.005 | 0.708±0.003 0.676±0.037 0.777±0.025 | 8 7-10 5-6 | 1.490±0.025 1.545±0.120 1.947±0.093 | 0.835±0.030 0.763±0.034 0.710±0.039 |
| Network | Alpha () | AUC | Accuracy | Live Acc. | Spoof Acc. | Placement | d’ | FNR@FPR=1% |
| ResNet DenseNet Inception | 0.9 | 0.914±0.007 0.940±0.001 0.953±0.004 | 0.830±0.005 0.860±0.010 0.868±0.006 | 0.937±0.007 0.951±0.010 0.961±0.005 | 0.714±0.006 0.762±0.031 0.769±0.016 | 5 5 4-5 | 1.932±0.061 2.198±0.015 2.376±0.054 | 0.576±0.025 0.466±0.023 0.413±0.039 |
| ResNet DenseNet Inception | 0.7 | 0.929±0.009 0.937±0.005 0.948±0.005 | 0.845±0.014 0.853±0.011 0.860±0.005 | 0.933±0.026 0.941±0.011 0.952±0.002 | 0.750±0.021 0.759±0.029 0.761±0.012 | 5 5 5 | 2.080±0.097 2.169±0.055 2.296±0.060 | 0.582±0.065 0.551±0.051 0.472±0.030 |
| ResNet DenseNet Inception | 0.5 | 0.918±0.007 0.930±0.008 0.945±0.004 | 0.837±0.008 0.842±0.011 0.860±0.003 | 0.930±0.004 0.941±0.005 0.942±0.006 | 0.736±0.014 0.734±0.020 0.772±0.008 | 5 5 5 | 1.972±0.062 2.088±0.082 2.258±0.055 | 0.602±0.044 0.559±0.035 0.538±0.062 |
| ResNet DenseNet Inception | 0.3 | 0.920±0.001 0.918±0.009 0.934±0.009 | 0.843±0.004 0.824±0.010 0.845±0.013 | 0.915±0.014 0.935±0.002 0.923±0.018 | 0.765±0.013 0.703±0.022 0.761±0.026 | 5 5-7 5 | 1.991±0.014 1.969±0.082 2.131±0.102 | 0.615±0.033 0.651±0.025 0.618±0.047 |
| ResNet DenseNet Inception | 0.1 | 0.878±0.011 0.887±0.015 0.900±0.001 | 0.803±0.011 0.800±0.015 0.819±0.004 | 0.850±0.011 0.886±0.009 0.873±0.015 | 0.752±0.033 0.706±0.021 0.760±0.025 | 6-7 7-8 5-6 | 1.654±0.082 1.721±0.111 1.810±0.009 | 0.748±0.016 0.755±0.064 0.752±0.027 |
| Network | Alpha () | AUC | Accuracy | Live Acc. | Spoof Acc. | Placement | d’ | FNR@FPR=1% |
| ResNet DenseNet Inception | 0.9 | 0.924±0.014 0.942±0.012 0.955±0.002 | 0.835±0.017 0.861±0.008 0.875±0.003 | 0.925±0.008 0.948±0.009 0.965±0.002 | 0.738±0.032 0.768±0.011 0.777±0.007 | 5-7 5 4-5 | 2.031±0.132 2.236±0.148 2.396±0.028 | 0.593±0.040 0.468±0.050 0.387±0.011 |
| ResNet DenseNet Inception | 0.7 | 0.925±0.003 0.922±0.005 0.961±0.004 | 0.847±0.007 0.828±0.004 0.885±0.004 | 0.923±0.006 0.934±0.012 0.952±0.006 | 0.765±0.012 0.713±0.021 0.812±0.014 | 5 5-6 4 | 2.038±0.033 2.007±0.046 2.490±0.067 | 0.567±0.034 0.624±0.036 0.386±0.041 |
| ResNet DenseNet Inception | 0.5 | 0.924±0.014 0.926±0.007 0.959±0.002 | 0.853±0.011 0.843±0.009 0.880±0.002 | 0.928±0.012 0.926±0.015 0.950±0.009 | 0.771±0.035 0.753±0.004 0.806±0.011 | 5 5 4 | 2.031±0.137 2.050±0.075 2.453±0.034 | 0.479±0.010 0.546±0.050 0.416±0.028 |
| ResNet DenseNet Inception | 0.3 | 0.911±0.005 0.922±0.006 0.944±0.001 | 0.837±0.005 0.844±0.006 0.864±0.005 | 0.905±0.013 0.928±0.008 0.945±0.005 | 0.764±0.009 0.753±0.014 0.777±0.013 | 5 5 5 | 1.907±0.046 2.004±0.060 2.249±0.018 | 0.622±0.011 0.575±0.020 0.520±0.022 |
| ResNet DenseNet Inception | 0.1 | 0.878±0.003 0.896±0.009 0.932±0.001 | 0.809±0.004 0.823±0.010 0.852±0.006 | 0.857±0.009 0.891±0.023 0.907±0.001 | 0.756±0.019 0.750±0.027 0.794±0.012 | 7 5-7 5 | 1.646±0.021 1.784±0.070 2.106±0.014 | 0.764±0.016 0.694±0.017 0.631±0.023 |
| Network | Alpha () | AUC | Accuracy | Live Acc. | Spoof Acc. | Placement | d’ | FNR@FPR=1% |
| ResNet DenseNet Inception | 0.9 | 0.930±0.014 0.933±0.005 0.956±0.007 | 0.853±0.022 0.839±0.023 0.881±0.013 | 0.927±0.006 0.946±0.010 0.957±0.001 | 0.774±0.048 0.724±0.057 0.800±0.027 | 5-6 5-7 4-5 | 2.097±0.141 2.120±0.053 2.427±0.108 | 0.520±0.047 0.530±0.018 0.416±0.102 |
| ResNet DenseNet Inception | 0.7 | 0.918±0.003 0.930±0.006 0.943±0.002 | 0.835±0.011 0.846±0.005 0.860±0.005 | 0.922±0.014 0.939±0.006 0.951±0.003 | 0.740±0.020 0.745±0.013 0.761±0.013 | 5-6 5 5 | 1.968±0.032 2.088±0.064 2.234±0.022 | 0.593±0.061 0.557±0.051 0.485±0.046 |
| ResNet DenseNet Inception | 0.5 | 0.923±0.005 0.914±0.007 0.953±0.006 | 0.842±0.010 0.830±0.011 0.871±0.006 | 0.918±0.010 0.922±0.009 0.949±0.009 | 0.759±0.016 0.731±0.032 0.786±0.007 | 5 5-6 4-5 | 2.022±0.052 1.933±0.066 2.374±0.088 | 0.550±0.042 0.658±0.043 0.447±0.016 |
| ResNet DenseNet Inception | 0.3 | 0.913±0.015 0.909±0.008 0.940±0.007 | 0.838±0.014 0.838±0.004 0.861±0.010 | 0.902±0.016 0.905±0.016 0.934±0.002 | 0.769±0.015 0.767±0.010 0.783±0.018 | 5-6 5 5 | 1.933±0.126 1.893±0.071 2.198±0.081 | 0.600±0.022 0.681±0.056 0.521±0.043 |
| ResNet DenseNet Inception | 0.1 | 0.886±0.007 0.893±0.008 0.924±0.005 | 0.815±0.008 0.819±0.006 0.846±0.004 | 0.858±0.012 0.895±0.014 0.901±0.011 | 0.768±0.016 0.736±0.010 0.787±0.010 | 5-7 5-7 5 | 1.704±0.054 1.756±0.064 2.026±0.051 | 0.759±0.009 0.745±0.053 0.665±0.046 |
| Network | Alpha () | AUC | Accuracy | Live Acc. | Spoof Acc. | Placement | d’ | FNR@FPR=1% |
| ResNet DenseNet Inception | 0.9 | 0.912±0.006 0.929±0.004 0.951±0.003 | 0.828±0.012 0.841±0.006 0.867±0.001 | 0.925±0.009 0.945±0.014 0.961±0.008 | 0.724±0.015 0.728±0.020 0.765±0.009 | 5-6 5 5 | 1.912±0.054 2.079±0.042 2.338±0.046 | 0.576±0.049 0.509±0.024 0.430±0.037 |
| ResNet DenseNet Inception | 0.7 | 0.909±0.016 0.919±0.008 0.953±0.006 | 0.825±0.022 0.823±0.005 0.871±0.010 | 0.929±0.003 0.928±0.011 0.950±0.008 | 0.712±0.043 0.710±0.008 0.785±0.027 | 5-7 5-6 4-5 | 1.898±0.135 1.976±0.074 2.373±0.083 | 0.576±0.044 0.643±0.036 0.463±0.029 |
| ResNet DenseNet Inception | 0.5 | 0.914±0.008 0.902±0.005 0.940±0.007 | 0.830±0.009 0.800±0.004 0.860±0.003 | 0.935±0.002 0.931±0.006 0.936±0.012 | 0.717±0.018 0.659±0.011 0.779±0.008 | 5-6 7 5 | 1.934±0.069 1.826±0.041 2.199±0.090 | 0.613±0.011 0.683±0.061 0.503±0.045 |
| ResNet DenseNet Inception | 0.3 | 0.903±0.007 0.869±0.015 0.922±0.012 | 0.826±0.005 0.777±0.008 0.832±0.016 | 0.892±0.016 0.899±0.015 0.931±0.008 | 0.755±0.027 0.645±0.014 0.726±0.027 | 5-6 7-10 5-6 | 1.841±0.056 1.593±0.103 2.015±0.126 | 0.669±0.047 0.740±0.035 0.602±0.027 |
| ResNet DenseNet Inception | 0.1 | 0.844±0.035 0.785±0.017 0.756±0.012 | 0.772±0.030 0.708±0.010 0.711±0.008 | 0.874±0.018 0.819±0.017 0.808±0.012 | 0.662±0.045 0.589±0.020 0.607±0.004 | 7-11 12 12 | 1.444±0.211 1.118±0.082 0.982±0.055 | 0.766±0.028 0.839±0.020 0.780±0.044 |
| Network | Alpha () | AUC | Accuracy | Live Acc. | Spoof Acc. | Placement | d’ | FNR@FPR=1% |
| ResNet DenseNet Inception | 0.9 | 0.931±0.002 0.931±0.006 0.955±0.004 | 0.836±0.010 0.846±0.013 0.873±0.005 | 0.941±0.004 0.940±0.009 0.957±0.010 | 0.722±0.019 0.745±0.036 0.782±0.009 | 5-6 5 4-5 | 2.093±0.017 2.102±0.067 2.404±0.068 | 0.582±0.016 0.529±0.024 0.441±0.078 |
| ResNet DenseNet Inception | 0.7 | 0.911±0.018 0.936±0.004 0.951±0.005 | 0.831±0.014 0.858±0.002 0.874±0.007 | 0.921±0.016 0.932±0.003 0.948±0.004 | 0.734±0.019 0.777±0.006 0.795±0.019 | 5-7 5 4-5 | 1.920±0.155 2.156±0.044 2.350±0.077 | 0.614±0.062 0.540±0.025 0.469±0.029 |
| ResNet DenseNet Inception | 0.5 | 0.907±0.014 0.911±0.009 0.953±0.007 | 0.830±0.018 0.828±0.012 0.876±0.011 | 0.910±0.023 0.902±0.014 0.953±0.014 | 0.743±0.023 0.748±0.015 0.793±0.023 | 5-7 5-7 4-5 | 1.872±0.116 1.903±0.074 2.378±0.097 | 0.678±0.092 0.662±0.036 0.424±0.051 |
| ResNet DenseNet Inception | 0.3 | 0.882±0.002 0.902±0.003 0.932±0.000 | 0.810±0.006 0.820±0.005 0.852±0.006 | 0.889±0.006 0.911±0.003 0.928±0.010 | 0.725±0.006 0.722±0.007 0.770±0.023 | 6-7 5-6 5 | 1.674±0.016 1.831±0.028 2.111±0.003 | 0.765±0.053 0.703±0.021 0.635±0.010 |
| ResNet DenseNet Inception | 0.1 | 0.860±0.001 0.872±0.006 0.899±0.008 | 0.786±0.001 0.783±0.005 0.817±0.007 | 0.853±0.004 0.869±0.006 0.876±0.011 | 0.713±0.005 0.691±0.009 0.754±0.011 | 7 7-8 5-7 | 1.531±0.009 1.605±0.040 1.806±0.062 | 0.873±0.026 0.801±0.021 0.751±0.042 |

### v2

@@ 1 Introduction
Alongside contributions at the intersection of saliency-guided training and fingerprint PAD, this paper releases all explored saliency data at https://github.com/CVRL/Human-Machine-
Pairing-Fingerprint-PAD, including the 800 doubly human-annotated saliency maps, the various sets of algorithmically-sourced pseudosaliency and the trained fingerprint PAD models.
@@ 3.1 Justification of Data Selection
We use the LivDet-Fingerprint datasets from 2015, 2017, 2019, and 2021 for both saliency collection and saliency-guided training [20, 21, 25, 5]. These datasets were chosen for three key reasons. First, they are the most recent and well-established benchmarks for evaluation of fingerprint PAD available to these authors. Second, their combined use provides a broader range of presentation attack instruments and capture devices, providing greater diversity than any single LivDet-Fingerprint competition alone. Finally, the LivDet-Fingerprint 2021 test set is the most challenging among our available datasets, as evidenced by its lower reported competitor scores compared to previous available LivDet-Fingerprint test sets. This increased difficulty, along with its inclusion of attack types not present in its training set or any prior LivDet dataset, makes it a strong choice for evaluating model performance.
@@ 3.2 Construction of Limited Data Context
The training of biometric PAD models, including fingerprint PAD, is often limited by the availability of high-quality spoof samples [17]. To simulate training conditions under a limited data availability context, we construct an 800-sample training dataset. This dataset size is selected to align with similar studies exploring low-data saliency strategies (e.g. 765 in [2]). The limited data context is balanced by class, with 400 bonafide and 400 spoof samples. The spoof samples are further balanced over eight attack types at 50 samples each, sourced from the LivDet-Fingerprint 2015-2019 train and test datasets and the LivDet-Fingerprint 2021 train datasets: Ecoflex [20, 21, 25], gelatine [20, 21, 25], latex [20, 21, 25, 5], liquid Ecoflex [20, 21, 25], woodglue [20, 21, 25], body double [21], mix [25], and RPRO Fast [5]. Bonafide and spoof samples are further balanced by the capture sensors represented by each subclass (note that not all attack types are represented by all scanners).
@@ 3.3 Human-Annotative Saliency Collection
We recruited 50 participants from the University of Notre Dame using a convenience sampling technique due to accessibility, scheduling, and other practical constraints. We collected annotations over the limited 800-sample dataset described in Sec. 3.2. Each participant annotated 32 fingerprint scans, 16 bonafide and 16 spoof samples, with spoof samples spanning eight attack types (two images per type). Participants used an image annotation software specially designed for this study to (a) predict each sample’s class (genuine or fake), (b) annotate salient regions supporting their decision, and (c) optionally provide textual description to further support their classification decisions (e.g., in cases when it would be difficult to support subject’s decision by annotating the image, such as perfect symmetry of the fingerprint image, or unnatural contrast, or too dark image). Participants were not instructed to annotate any specific fingerprint features (e.g. whorls, ridges), as intuitively-annotated regions mitigate feature bias, promote straightforward collection reproducibility, and better align with the philosophy of saliency guidance. The created image annotation software also recorded start and end time stamps of each stroke. This allows for fine-grained analysis of annotation time and provides the order of features annotated by each subject on each sample. Confounding factors in this study include the level of biometric expertise or fingerprint familiarity among annotators. However, we consider this factor to be of limited concern, as prior saliency collection studies have demonstrated success with non-expert annotators [2, 3].
@@ 5 Results
Among human-saliency configurations with limited-data loss-based guidance (scenario S3), the configuration achieving the highest accuracy is configured with the Area of Interest saliency granularity, the Inception architecture, and a CYBORG weight of 0.9. Its achieved accuracy of 0.879±0.006 and AUC of 0.959±0.002 outperform the best baseline configuration (scenario S2) by 0.017 and 0.013, respectively.
@@
The generative nature of pseudosaliency enables saliency-guided training to scale effectively in larger data availability contexts, as explored in scenario S5. The best configuration in this scenario employs minutiae-based Area of Interest pseudosaliency, achieving an accuracy of 0.938±0.004. This result outperforms the large-data baseline (scenario S1) by 0.027. When compared to the reported LivDet-Fingerprint 2021 rankings [5], this performance would have earned or tied for first place, accounting for our reported standard deviation (which was not included in the LivDet-Fingerprint 2021 results). This configuration raises spoof detection accuracy by 8.9%, demonstrating the generalization strength of saliency-guided training.
@@ 6 Discussion
A closer look at model performance metrics in Tab. 1 reveals that saliency-guided training primarily raises spoof detection accuracy, with bonafide accuracy usually affected to a lesser degree. In the limited-data context, the best configuration (Inception, CYBORG, =0.7, Minutiae FOI) improves spoof accuracy by 4.3%, while bonafide accuracy increases by only 0.4%. Similarly, in the large-data context, the best configuration (DenseNet, CYBORG, =0.9, Minutiae AOI) gains 8.9% in spoof accuracy but loses 3.1% in bonafide accuracy.
@@ 6.3 Blurring and BOI Saliency Granularity
In blur-based saliency guidance, the Boundary of Interest (BOI) saliency granularity consistently performs well. It is the best granularity for human saliency and second best among explored pseudosaliency types. BOI saliency identifies a minimal enclosing rectangle around all regions of interest, which typically encompasses most of the fingerprint area. Consequently, when non-salient regions are blurred, the fingerprint pattern remains largely unaffected. Combined with eightfold data augmentation through multiple blur radii, the model encounters the nearly unblurred pattern repeatedly. This inadvertently increases exposure to samples during training, similar to training with standard cross entropy for many additional epochs. While BOI saliency achieves high scores on key metrics, this effectiveness is likely due to unintended augmentation behavior rather than true saliency-based guidance, which impacts how blur-based saliency guidance results may be interpreted.
@@ 7 Conclusion
Acknowledgment
This material is based upon work supported by the U.S. National Science Foundation under grant No. 2237880. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. National Science Foundation.
@@ Acknowledgment
Accuracy: Overall classification accuracy on the LivDet-Fingerprint 2021 test set.
Bonafide Accuracy: Classification accuracy on bonafide (live, real) fingerprint samples from the LivDet-Fingerprint 2021 test set.
Spoof Accuracy: Classification accuracy on spoof (fake) fingerprint samples from the LivDet-Fingerprint 2021 test set.
@@
Accuracy: Classification accuracy across the various spoof (fake) types in the LivDet-Fingerprint 2021 test set. A superscript indicates an unknown attack type (no samples of this attack type seen during training) while a superscript indicates a known attack type (samples of this attack type seen during training).
Appendix A Scenario S1, Full Results
| Network | AUC | Accuracy |
Placement | d’ |
@@ Appendix B Scenario S2, Full Results
| Network | AUC | Accuracy |
Placement | d’ |
@@ Appendix C Scenario S3, Full Results
| Network | Alpha () | AUC | Accuracy |
Placement | d’ |
0.9 |
0.7 |
0.5 |
0.3 |
0.1 |
| Network | Alpha () | AUC | Accuracy |
Placement | d’ |
0.9 |
0.7 |
0.5 |
0.3 |
0.1 |
| Network | Alpha () | AUC | Accuracy |
Placement | d’ |
0.9 |
0.7 |
0.5 |
0.3 |
0.1 |
| Network | Alpha () | AUC | Accuracy |
Placement | d’ |
0.9 |
0.7 |
0.5 |
0.3 |
0.1 |
| Network | Alpha () | AUC | Accuracy |
Placement | d’ |
0.9 |
0.7 |
0.5 |
0.3 |
0.1 |
| Network | Alpha () | AUC | Accuracy |
Placement | d’ |
0.9 |
0.7 |
0.5 |
0.3 |
0.1 |
| Network | Alpha () | AUC | Accuracy |
Placement | d’ |
0.9 |
0.7 |
0.5 |
0.3 |
0.1 |